Let X∼Poisson(k−1)X∼Poisson(k−1) with Var(X)=1Var(X)=1, then find the value of the constant kk


1
Expert's answer
2021-04-29T16:27:37-0400

We know that if "X" follows a Poisson distribution of parameter "\lambda" , i.e. "X\sim \textrm{Poisson}(\lambda)" , then the variance of "X" is "\textrm{Var}(X)=\lambda".

If we apply it to our case we have:


"\textrm{Var}(X)=k-1=1\Rightarrow k-1=1\Rightarrow k=1+1\Rightarrow k=2"


Thus, the value of the constant "k" is "k=2" .